OpenStudy (anonymous):

how to check 17=p-3-3p if -10=p

OpenStudy (kaiz122):

just substitute -10 to the p in the equation.

OpenStudy (shaik0124):

just substitute p=10 in 17=p-3-3p u will lefthand and right side equal

OpenStudy (shaik0124):

sorry p=-10

OpenStudy (anonymous):

but how do you check it when i try i get 17=(-10)-3-3(-10) 17=7-(-30) am i correct? after i get there i get confused

OpenStudy (kaiz122):

of both sides are equal.

OpenStudy (shaik0124):

see u consider the equation in two parts that is left hand side as 17 and right hand side as 17=p-3-3p now put p=-10 in rhside that will be equal to lhside which is hence proved lhs=rhs

OpenStudy (shaik0124):

@yanna did u got my point

OpenStudy (anonymous):

no because when i did 7- -30 i got -23 and that is not equal to 17

OpenStudy (shaik0124):

i will show the work wait@yana

OpenStudy (kaiz122):

i thinl it should be -13 --30

OpenStudy (kaiz122):

-10-3=-13

OpenStudy (kaiz122):

-13+30=17

OpenStudy (anonymous):

yes that makes a little more sense to me thank @kaiz122

OpenStudy (shaik0124):



OpenStudy (shaik0124):

@yana